A series R-L circuit is connected to 110V ac source.If voltage across resistor is 85V ,find power dissipated in circuit


1
Expert's answer
2021-05-31T15:15:46-0400

Since the voltages across the inductor and the resistor are 90° out of phase, we obtain the voltage across the inductor in the phase domain as follows: If the voltage across the resistor is V1

The voltage across the inductor will be jV2


. Assume the voltage of the source is V3

Applying kvL law


"V_1+jV_2=V_1cos(wt)+V_2sin(wt)=V_3cos(wt-\theta)=85+V_2sin(wt)=110cos(wt-\theta)"

"110=\sqrt{85^2+V_2^2}"

"V_2=69.82V"

For example

"P_{ave}=\frac{1}{2}\frac{V_R^2}{R}"

Assume

"R=20 ohm"

And VR=85 V

Put value

"P_{av}=\frac{1}{2}\frac{85^2}{20}=180.6 W"